OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____. 3x= 2 3x= 5 5x= 5

OpenStudy (anonymous):

plug x = -1 + 2y into 2x + y = 3 then you'll get 2(-1 + 2y) + y = 3 >>> -2 +4y +y = 3 solve for y, and substitute into x = -1 + 2y and solve for x

OpenStudy (anonymous):

5X=5